# The Superconducting Quasiparticle-Amplifying Transmon: A Qubit-Based Sensor for meV Scale Phonons and Single THz Photons

### v3

@@ Appendix A Resonator Readout Modes
When measuring the qubits, we can operate in two distinct modes: amplitude or phase readout. Recall that our voltage signal is the transmission signal of a complex pulsed readout tone $V_{r}$
@@
Amplitude Readout—Since $S_{21}$ can only take on values between zero and one, the maximum and minimum amplitudes of the signal voltage are $|V_{s}|=|V_{r}|$ and $|V_{s}|=0$ respectively. Optimal amplitude readout is accomplished in practice by designing a state separation $2\chi$ that is much larger than the bandwidth of the readout ($2\chi\gg f_{0}/Q$). Thus the readout frequency can be set at one parity state while the other state is outside the bandwidth of the readout. This can be seen visually on the IQ circle in Fig. 5.
Phase Readout—The opposite readout technique would be to consider only the phase of the signal tone. The maximum phase separation between the two states occurs where the amplitude of the readout signal is the same, or rather
@@
The optimization problem becomes finding the $\chi$ to maximize the phase separation between $S_{21,e}$ and $S_{21,o}$. Let us first recall the definition of $S_{21}$ when limited by the quality factor of the coupling
@@
recalling that the effective dispersion $\chi=\chi_{0}\cos(\pi n_{g})$.
For simplicity, let us write $S_{21}$ as
@@
Considering two points on the IQ unit circle (see Fig. 5), the maximum separation (maximum signal) of phase between the points occurs when they are separated by a phase shift of $\pi/2$. Since the signal is symmetric about the I axis, the maximum corresponds to $\tan(\pm\pi/4)=\pm 1$.
@@ Appendix B Readout Fidelity
Given the simplified readout modes described in Appendix A, we are prepared now to talk about readout fidelity for this type of sensor design. We will show, as is typical of readout in I,Q space, that proper optimization of the design of the sensor will reduce sensitivity to readout frequency within a range of the nominal central frequency determined by $Q$. For those familiar with this general optimization, proceed to Equation 43 for the discussion of how photon saturation affects readout optimization.
@@
Let’s now consider the general case for the optimization of the phase readout (in which we assume that we have access to both the phase and amplitude of $S_{21}$). Our optimization ensures that the even and odd states lie opposite each other on the resonance circle. For a rotation angle of $\theta$ about the point (I,Q) = (1/2,0), we find that
@@
Thus the design optimized for phase readout will have the same signal amplitude regardless of the readout frequency, as the two points will be shifted an equal amount around the resonator circle but maintain unit distance between them. Contrast this with amplitude readout in the non-optimized case, where maximum signal is only achieved when the readout frequency is exactly the frequency of one of the parity states.
@@
$\epsilon_{r}=1-\exp\left(-\frac{\hbar f_{0}^{2}}{Q_{c}P_{r}}\right)$ (43)
The frequency dependence comes both from calculating the decay rate for a given coupling factor $Q_{c}$ and from converting readout power into photon number. The readout efficiency improves for lower Q or higher readout frequency; both of these effects allow the SQUAT to be interrogated ‘more often’ because the absorbed photons in the SQUAT decay back into the feedline more rapidly, allowing another photon from the feedline to be absorbed. Eq. 43 also shows that for high readout power we suffer from the limited number of photons interacting with the SQUAT.
Having established that our signal amplitude is always equal to our readout power, we now consider the noise contribution to our signal in order to determine readout fidelity. Assuming we are limited by Johnson-Nyquist noise (which is uncorrelated to the signal), we find a noise power in an impedance-matched network of
@@ B.1 Sources of Phase Noise
In the case of charge noise, we can use the gate capacitance to estimate the effect of a known gate voltage noise on the system; TLS noise will be determined by the occupation and coupling of, e.g., oxide states near the qubit to the qubit islands.
@@ Appendix C Quasiparticle Diffusion Model
This diffusion model describes the dynamics of a non-equilibrium QP population in the qubit islands and how they are measured. As described in the text above, this QP population arises from an impinging particle, e.g. a phonon or photon, with energy greater than twice the superconducting bandgap of the island. The incident particle breaks a Cooper-pair into highly energetic QPs, which then undergo a downconversion process into phonons and lower energy QPs until a quasi-stable population of QPs exists at the superconducting band edge [19]. The following model begins after this downconversion step. A schematic diagram of the full process can be seen in Fig. 1.
@@ C.1 QP Diffusion in Island
$\frac{\partial}{\partial t}n(\mathbf{x},t)=D_{\text{island}}\nabla^{2}n(\mathbf{x},t)-\frac{n(\mathbf{x},t)}{\tau_{\text{island}}}+s\,,$ (49)
@@
where $n(\mathbf{x},t)$ is the QP number density, $D_{\text{island}}$ and $\tau_{\text{island}}$ are the diffusion coefficient and finite QP lifetime in the island (Al in this case), and $s$ is a source of particles. This differential equation can be solved in two dimensions; from [71], the fraction of QPs that get trapped in the overlapping region is given by
$F_{c}^{2D}=\frac{2\rho_{1}}{\rho_{2}^{2}-\rho_{1}^{2}}\frac{I_{1}(\rho_{2})K_{1}(\rho_{1})-I_{1}(\rho_{1})K_{1}(\rho_{2})}{I_{1}(\rho_{2})\left[K_{0}(\rho_{1})+\lambda_{a}K_{1}(\rho_{1})\right]+K_{1}(\rho_{2})\left[I_{0}(\rho_{1})-\lambda_{a}I_{1}(\rho_{1})\right]},$ (50)
where $I_{i}$ and $K_{i}$ are modified Bessel functions of the first and second kind and
$\rho_{1}\equiv\frac{R_{1}}{L_{d}},\quad\rho_{2}\equiv\frac{R_{2}}{L_{d}},\quad\lambda_{a}\equiv\frac{L_{a}}{L_{d}}.$ (51)
$R_{1}$ and $R_{2}$ are the effective radii of the island and trapping regions respectively, and $L_{d}$ and $L_{a}$ are two competing length scales: the diffusion length and absorption length, given by
$L_{d}\equiv\sqrt{D_{\text{island}}\tau_{\text{island}}}\approx\alpha_{\text{island}}h_{\text{island}}$ (52)
$L_{a}\equiv\frac{D_{\text{island}}}{\nu_{abs}}\approx\frac{1}{p_{abs}}\frac{h_{\text{island}}^{2}}{\ell_{\text{trap}}},$ (53)
@@
where $\alpha_{\text{island}}$ is a film material-specific (Al in this work) scalar for the case that the diffusion length is limited by the film thickness, $h_{\text{island}}$ is the island thickness, $\ell_{\text{trap}}$ is the length of the overlapping region between the two materials, we refer to $\nu_{abs}$ as the effective QP absorption velocity, and $p_{abs}$ is the per QP transmission probability at the interface.
@@ C.2 Quasiparticle Multiplication
When the QPs enter the trap they will have an energy of roughly $\Delta_{\text{Island}}$ (see Fig. 1 b). Once in the lower $T_{c}$ trap however, these QPs will undergo a downconversion step into a lower energy QP population and athermal phonons. During the down-conversion process, a fraction of the athermal phonons created have less energy than $2\Delta_{\text{Trap}}$ and will be lost into the substrate. The energy that remains in the QP system ($\eta_{\text{QP}}$) depends on the incident energy of the original particle, in this case of energy $\Delta_{\text{Island}}$. If the initial QP has energy of $2\Delta_{\text{Trap}}$, then $\eta_{\text{QP}}$ is approximately unity. This is because the QPs are already roughly at the band edge and will not undergo downconversion. If the incident particle has energy greater than $2\Delta_{\text{Trap}}$ or less than $4\Delta_{\text{Trap}}$, $\eta_{\text{QP}}$ decreases significantly, because the only kinimatically allowed phonons released will be sub-gap, and thus not contribute to the QP population, roughly half the energy will go into sub-gap phonons. For incident energies above this range, $\eta_{\text{QP}}$ tends to asymptote to a fixed value, this value depends in part on QP lifetimes and above-gap phonon loss mechanisms that are material dependent. A simple simulation, ignoring QP and phonon lifetime effects, of the energy in the QP system as a function of incident particle energy, in this case $\Delta_{\text{Island}}/\Delta_{\text{Trap}}$, is shown Fig. 8.
@@
which, as shown in Fig. 8, when the ratio becomes greater than $\sim 3$, $\eta_{\text{QP}}$ starts to become constant at a value of roughly $60\%$, and the gain scales linearly with the ratio. For our targeted island-to-trap gap ratio of $\sim 10$, we should expect a QP number gain of about 7.
Figure 8: Left Axis: Simulated percentage of energy remaining in the QP system ($\eta_{\text{QP}}$) as a function of incident particle energy relative to the trap gap. In this case the incident particle is a QP of energy $\Delta_{\text{Island}}$, in units of the trapping material’s gap $\Delta_{\text{Trap}}$. Right Axis: Simulated QP number gain, Eq. 56, as a function of SC gap ratio using $\eta_{\text{QP}}$ from the same figure.
@@ C.3 Quasiparticle Diffusion in Trap
At this point, a stable population of QPs has developed in the trap. We then model the trap QP transport using a 1D diffusion model, which is justified in this case since the length of this region is much larger than the width or film thickness. Solving Eq. 49 in one dimension, the fraction of QPs tunneling across the junction is given by
@@
where $\ell_{\text{JJ}}$ is the length of the Josephson junction, $\alpha_{\text{trap}}$ is a scalar quantifying the diffusion in the thin film thickness limited case, and $p_{\text{tunnel}}$ is the per-QP tunneling probability at the junction. Since the relative amount of Mn in the AlMn should be very low, the QP scattering diffusion length should be close to that of a typical Al thin film. As such, we use a value of $\alpha_{\text{AlMn}}\approx 570$ as measured in [76]. From [77] we estimate the QP tunneling probability to be $p_{tunnel}\approx 1-10\,(\times 10^{-6})$ for pure Al. For the AlMn in this model, we adopt the conservative lower bound of $1\times 10^{-6}$.

### v4

@@ Appendix A Resonator Readout Modes
When measuring time-domain SQUAT behavior in a highly multiplexed environment, we are unlikely to be able to retain both quadrature components, and we want to see how streaming only a single quadrature will impact our ability to discriminate between parity states. This is not the optimal readout mode, as retaining both quadratures allows for optimal signal to noise as discussed in the next appendix. This appendix serves instead to illustrate how we should optimize dispersion such that these two limited cases can have maximal sensitivity to absorbed energy.
We can operate SQUATs in two distinct modes: amplitude or phase readout. For this discussion, recall that our voltage signal is the transmission signal of a complex pulsed readout tone $V_{r}$:
@@
Amplitude Readout—Since $|S_{21}|$ can only take on values between zero and one, the maximum and minimum amplitudes of the signal voltage are $|V_{s}|=|V_{r}|$ and $|V_{s}|=0$ respectively. Optimal amplitude readout is accomplished in practice by designing a state separation $2\chi$ that is much larger than the bandwidth of the SQUAT ($2\chi\gg f_{0}/Q$). Thus the readout frequency can be set at one parity state while the other state is outside the bandwidth of the readout. This can be seen visually on the IQ circle in Fig. 6. This is the readout mode we would be forced to use if only one state lies on the resonator circle, which is the case for the condition $2\chi\gg f_{0}/Q$. We would only have maximal contrast between states for such a device with a tone placed directly on resonance with the SQUAT, which would represent maximal energy coupling into the device.
Phase Readout—Consider instead if we were to optimize the SQUAT dispersion such that we can maximize state contrast in the phase of the transmitted signal tone. Microphyiscally, this corresponds to photons scattering off of the SQUAT and picking up a state-dependent phase shift. To solve for the optimal dispersion given a known SQUAT quality factor, we first simplify the problem such that we remove all information from the amplitude readout axis. This corresponds to the condition
@@
Solving this equality thus gives us an optimization in which both amplitude and phase readouts contain maximal parity information. If this is the case, as we will show in the next appendix, we can guarantee that full quadrature readout at any frequency between the two parity states has an equal amount of discrimination, which will be important for optimizing sensor fidelity.
Let us first recall the definition of $S_{21}$ when limited by the quality factor of the coupling
@@
recalling that the effective dispersion $\chi=\chi_{0}\cos(\pi n_{g})$. For simplicity, let us write $S_{21}$ as
@@
Considering two points on the IQ unit circle (see Fig. 6), the maximum separation (maximum signal) of phase between the points occurs when they are separated by a phase shift of $\pi/2$. Any solution with this phase shift will be valid, but this yields the simplest solution in the case of the phase readout described earlier, where the states are symmetric about the I axis.
@@
For I symmetric states, the maximum separation corresponds to $\tan(\pm\pi/4)=\pm 1$.
@@
which is essentially a condition for the minimum dispersion we need to design into the SQUAT to ensure we can always tune the charge bias to find an optimally separated set of parity states. In the case where we expect significant charge noise, there is some value in designing a more fine-tuned dispersion as close to this value as possible; we will discuss this in the context of non-ideal noise sources in the next appendix.
As a final note, we point out that $Q_{c}$ is not itself an intrinsic quantity, but instead is defined as $Q_{c}=\frac{f_{0}}{\Gamma_{d}}$, where $\Gamma_{d}$ is the decay rate of photons from the SQUAT back into the environment. We thus find that this condition is just $2\chi=\Gamma_{d}$, or in other words, we want to ensure that the SQUAT decays quickly enough that we can distinguish between parity bands. This decay rate is also essentially the maximum interrogation rate of the SQUAT, and so tradeoffs in readout bandwidth dictate limits on the allowable dispersion of the SQUAT.
@@ Appendix B Readout Fidelity
Given the discussion of optimizing SQUAT dispersion vs bandwidth in Appendix A, we are prepared now to talk about readout fidelity for this type of sensor design. We will show, as is typical of readout in I,Q space, that proper optimization of the design of the sensor will reduce sensitivity to readout frequency within a range of the nominal central frequency determined by $Q$. For those familiar with this general optimization, proceed to Equation 43 for the discussion of how photon saturation affects readout optimization.
Signal Magnitude
@@
Let’s now consider the general case for the optimization of the phase readout (in which we assume that we have access to both the phase and amplitude of $S_{21}$). Tuning the SQUAT to a point where $2\chi\approx\frac{f_{0}}{Q_{c}}$ ensures, as described earlier, that the even and odd states lie opposite each other on the resonance circle. For a rotation angle of $\theta$ about the point (I,Q) = (1/2,0), we find that
@@
Thus the design tuned to this optimal dispersion will have the same signal amplitude regardless of the readout frequency, as the two points will be shifted an equal amount around the resonator circle but maintain unit distance between them. Contrast this with amplitude readout in the non-optimized case, where maximum signal is only achieved when the readout frequency is exactly the frequency of one of the parity states.
SQUAT Saturation
@@
$\epsilon_{r}=1-\exp\left(-\frac{\hbar f_{0}^{2}}{Q_{c}P_{r}}\right)=1-\exp\left(-\frac{\Gamma_{d}}{\Gamma_{\gamma}}\right)$ (43)
The frequency dependence comes both from calculating the decay rate for a given coupling factor $Q_{c}$ and from converting readout power into photon number.
If we convert power to photon interrogation rate $\Gamma_{\gamma}$ and substitute in our decay rate from earlier (second half of Eq.43), this becomes clearer. The readout efficiency improves when the interrogation rate is slower than the decay rate. Increasing the decay rate allows for the SQUAT to be readout more quickly, but comes with the side-effect of also coupling it more strongly to the environment. As discussed earlier, increasing decay rate also increases the minimum state dispersion, reducing the multiplexing factor. Finally, Eq. 43 also shows that for high readout power we suffer from the limited number of photons interacting with the SQUAT.
Readout Fidelity
Having established that our signal amplitude is always equal to our readout power times readout efficiency $P_{s}=\epsilon_{r}P_{r}$, we now consider the noise contribution to our signal in order to determine readout fidelity. Assuming we are limited by Johnson-Nyquist noise (which is uncorrelated to the signal), we find a noise power in an impedance-matched network of
@@ B.1 Sources of Phase Noise
In the case of charge noise, we can use the gate capacitance to estimate the effect of a known gate voltage noise on the system; TLS noise will be determined by the occupation and coupling of, e.g., oxide states near the qubit to the qubit islands. As mentioned earlier, the impact of charge noise is affected by the maximum qubit dispersion, as in the frequency-insensitive readout mode we will have a larger charge noise if the dispersion needs to be tuned down to readout the optimal readout point. This may mean that, in a high charge noise environment, we will have to resort to simpler magnitude readout and accept any resulting additional noise from frequency instability of the tone.
@@ Appendix C Quasiparticle Diffusion Model
This diffusion model describes the dynamics of a non-equilibrium QP population in the qubit islands and how they are measured. As described in the text above, this QP population arises from an impinging particle, e.g. a phonon or photon, with energy greater than twice the superconducting bandgap of the island. The incident particle breaks a Cooper pair into highly energetic QPs, which then undergo a downconversion process into phonons and lower energy QPs until a quasi-stable population of QPs exists at the superconducting band edge [19]. The following model begins after this downconversion step. A schematic diagram of the full process can be seen in Fig. 1.
@@ C.1 QP Diffusion in Island
$\frac{\partial}{\partial t}n(\mathbf{x},t)=D_{\text{island}}\nabla^{2}n(\mathbf{x},t)-\frac{n(\mathbf{x},t)}{\tau_{\text{island}}}+s(\mathbf{x},t)\,,$ (49)
@@
where $n(\mathbf{x},t)$ is the QP number density, $D_{\text{island}}$ and $\tau_{\text{island}}$ are the diffusion coefficient and finite QP lifetime in the island (Al in this case), and $s(\mathbf{x},t)$ is a source of particles.
Following the derivation from Moffatt [71], we assume that our source term $s(\mathbf{x},t)$ is a finite impulse, and we are interested in the total number of QPs absorbed. By integrating away the time degree of freedom, Eq. 49 becomes
$D_{\text{island}}\nabla^{2}\phi(\mathbf{x})-\frac{\phi(\mathbf{x})}{\tau_{\text{island}}}=-\sigma(\mathbf{x})$ (50)
where $\phi(\mathbf{x})$ is the scalar field defined as the time integrated particle number density and $\sigma(\mathbf{x})$ is the source density.
Figure 9: Schematic showing how the SQUAT sensor design (left) is mapped to two concentric circles (right)to model the QP diffusion and trapping process. Note that the scale of $R_{1}$ is enlarged in the figure simply for illustrative purposes.
We approximate the SQUAT as two concentric circles with outer radius $R_{2}=\ell_{\text{island}}$ and inner radius scales roughly as $R_{1}=\ell_{\text{trap}}$ as shown in Fig. 9. Since the film thickness will be $\mathcal{O}(10-100)\,\mathrm{nm}$, and the lengths of $\ell_{\text{trap}}$ and $\ell_{\text{island}}$ with be $\mathcal{O}(10-100)\,\mu\mathrm{m}$, we can treat the QP diffusion in the islands in two dimensions. Assuming that our QP source density is homogeneous in the island (a reasonable assumption for Al where the downconversion process is orders of magnitude shorter than the QP lifetime), we can write
$\sigma=\frac{N_{\text{stable},i}}{(R_{2}-R_{1})\pi},$ (51)
where $N_{\text{stable},i}$ is the total number of stable QPs after the initial downconversion process in the island. In cylindrical coordinates, Eq. 50 becomes
$r^{2}\frac{\partial^{2}}{\partial r^{2}}\phi+r\frac{\partial}{\partial r}\phi-\frac{1}{L_{d,i}^{2}}\phi=-\frac{\sigma(r)}{D_{\text{island}}}$ (52)
where $L_{d,i}\equiv\sqrt{D_{\text{island}}\tau_{\text{island}}}$ and the azimuthal angle has been suppressed due to cylindrical symmetry.
In our model, we assert that the outer edge of $R_{2}$ is a perfect reflector, and that we have some boundary that scales with $R_{1}$ that acts as a perfect absorber. Note that this perfect absorbing boundary does not correspond to the physical boundary of the trap, but rather we define the boundary with an effective characteristic absorption length scale (defined below) to simplify the math. From these assumptions we have the boundary conditions
$\frac{\partial}{\partial r}\phi(R_{2})=0$ (53)
$\frac{\partial}{\partial r}\phi(R_{1})L_{a,i}=\phi(R_{1}),$ (54)
where $L_{a,i}$ is defined as the absorption length, a weighting term to account for the non-perfect absorption at the trapping boundary. The absorption length can be thought of as the average distance a QP would have to travel in the trap in order to become ‘trapped’ to make the absorption equivalent to a step function boundary with an absorption probability less than unity. By weighting $R_{1}$ in this way, we can treat the boundary as a step function.
The differential equation in Eq. 52 can be solved along with the boundary conditions to be expressed in terms of modified Bessel functions [^1] In an effort to save space, only the final solution will be shown.. The total amount of QPs that are absorbed at the boundary defined by $R_{1}$ will be given by
$N_{\text{absorbed},i}=2\pi R_{1}D_{\text{island}}\frac{\partial}{\partial r}\phi(R_{1}).$ (55)
Finally we define the fraction of QPs that get trapped in the overlapping region vs the total number of stable QPs generated after the initial downconversion process in the island as
$F_{c}^{2D}=\frac{N_{\text{absorbed},i}}{{N_{\text{stable},i}}}=\frac{2\rho_{1}}{\rho_{2}^{2}-\rho_{1}^{2}}\frac{I_{1}(\rho_{2})K_{1}(\rho_{1})-I_{1}(\rho_{1})K_{1}(\rho_{2})}{I_{1}(\rho_{2})\left[K_{0}(\rho_{1})+\lambda_{a}K_{1}(\rho_{1})\right]+K_{1}(\rho_{2})\left[I_{0}(\rho_{1})-\lambda_{a}I_{1}(\rho_{1})\right]},$ (56)
where $I_{i}$ and $K_{i}$ are modified Bessel functions of the first and second kind. We note that while this derivation assumed the SQUAT was a cylinder, by the assumption of cylindrical symmetry and a homogeneous source term, this result is also valid for any wedge of the cylinder as well, e.g. each island of the SQUAT.
We now address the physicality of the assumptions of perfectly reflecting and absorbing boundaries by defining the diffusion and absorption length in terms of our physical design parameters. For a sufficiently pure Al film, the diffusion is limited by the film thickness and scales linearly with the thickness $h_{\text{island}}$
$L_{d,i}\equiv\sqrt{D_{\text{island}}\tau_{\text{island}}}\approx\alpha_{\text{island}}h_{\text{island}}$ (57)
@@
where $\alpha_{\text{island}}$ is a film material-specific scalar and is approximately 567 for aluminum [73]. Similarly, the absorption length depends on the thickness of the island on top of the trap, the length of the trap, and the probability of absorption. As shown in [18, 70], this is given by
$L_{a,i}\equiv\frac{D_{\text{island}}}{\nu_{abs}}\approx\frac{1}{p_{abs}}\frac{h_{\text{island}}^{2}}{\ell_{\text{trap}}},$ (58)
with $\ell_{\text{trap}}$ being the length of the overlapping region between the two materials, $\nu_{abs}$ the effective QP absorption velocity, and $p_{abs}$ the per QP transmission probability at the interface.
@@ C.2 Quasiparticle Multiplication
When the QPs enter the trap they will have an energy of roughly $\Delta_{\text{island}}$ (see Fig. 1 b). Once in the lower $T_{c}$ trap however, these QPs will undergo a downconversion step into a lower energy QP population and athermal phonons. During the downconversion process, a fraction of the athermal phonons created have less energy than $2\Delta_{\text{trap}}$ and will be lost into the substrate. The energy that remains in the QP system ($\eta_{\text{QP}}$) depends on the incident energy of the original particle, in this case of energy $\Delta_{\text{island}}$. If the initial QP has energy of $2\Delta_{\text{trap}}$, then $\eta_{\text{QP}}$ is approximately unity. This is because the QPs are already roughly at the band edge and will not undergo downconversion. If the incident particle has energy greater than $2\Delta_{\text{trap}}$ or less than $4\Delta_{\text{trap}}$, $\eta_{\text{QP}}$ decreases significantly, because the only kinimatically allowed phonons released will be sub-gap, and thus not contribute to the QP population, roughly half the energy will go into sub-gap phonons. For incident energies above this range, $\eta_{\text{QP}}$ tends to asymptote to a fixed value, this value depends in part on QP lifetimes and above-gap phonon loss mechanisms that are material dependent. A simple simulation of this is shown in Fig. 2. To simulate this, QP of energy $\Delta_{\text{island}}/\Delta_{\text{trap}}$ enters the trap and undergoes spontaneous downconversion, where a random percentage of its energy is given to a phonon and the rest remains with the QP, the above rules are then used to determine if the phonon is energetic enough break a cooper-pair into two new QPs. This process is repeated until all QPs are no longer energetic enough to generate pair-breaking phonons. In this simulation, it is assumed that both the QP and phonon lifetime is large compared with the scattering times.
@@
which, as shown in Fig. 2, when the ratio becomes greater than $\sim 3$, $\eta_{\text{QP}}$ starts to become constant at a value of roughly $60\%$, and the gain scales linearly with the ratio. For our targeted island-to-trap gap ratio of $\sim 10$, we should expect a QP number gain of about 7. We note that while this calculation assumes that all of the downconverted phonons that are energetic enough to generate new QPs will do so, we have the freedom in our design to dramatically increase the trapping thickness to minimize additional phonon losses and maximize the number of generated QPs. This value should be thought of as the upper bound for the maximum achievable gain.
@@ C.3 Quasiparticle Diffusion in Trap
At this point, a stable population of QPs has developed in the trap. We then model the trap QP transport using a 1D diffusion model, which is justified in this case since the length of this region is much larger than the width or film thickness. Since the QPs now have an energy much less than $2\Delta_{\text{island}}$, we treat the boundary between the edge of the trap and the island ($x=0$ in Fig. 10) as perfectly reflecting. We treat the junction as a perfect absorber (where we consider a ‘tunneled’ QP to be ‘absorbed’), and wrap the QP tunneling probability into the definition of the absorption length, similar to the two dimensional case above for the islands.
Figure 10: Schematic diagram of the trap and junction region (not to scale), showing the junction area, the trap, and the source term of QPs, $s(x,t)$.
In this case, we treat our source term in the diffusion equation as a delta function, both spatially and temporally, and Eq. 49 can be expressed in one dimension as
$\left(\partial_{x}^{2}-\frac{1}{L_{d,t}}^{2}\right)\phi=-\frac{N_{\text{stable, t}}}{D_{\text{trap}}}\delta(x-\ell_{\text{trap}}),$ (62)
Where again, $N_{\text{stable},t}$ is the total number of stable QPs after the downconversion/multiplication process in the trap. The above equation is solvable under the same boundary conditions as for the two dimensional case above, resulting in a fraction of tunneled QPs to the total number of stable QPs generated in the trap of
$F_{c}^{1D}=\frac{\lambda_{d}}{\lambda_{a}/\lambda_{d}+\coth\left(\frac{1}{\lambda_{d}}\right)},$ (63)
where $\lambda_{d}=L_{d,t}/\ell_{\text{trap}}$ and $\lambda_{a}=L_{a,t}/\ell_{\text{trap}}$.
Similar to before, we now account for the non-ideal absorption by relating the absorption and diffusion lengths to physical parameters. The diffusion in the trap will again be limited by surface effects and thus
$L_{d,t}\approx\alpha_{\text{trap}}h_{\text{trap}},$ (64)
where $h_{\text{trap}}$ is the trap thickness and $\alpha_{\text{trap}}$ is the scalar quantifying the diffusion in the thickness limited limit. We can describe the absorption the same way as for the two dimensional case, replacing absorption with tunneling effects,
$\lambda_{a}=\frac{1}{p_{\text{tunnel}}}\frac{h_{\text{trap}}^{2}}{\ell_{JJ}\ell_{\text{trap}}}$ (65)
where $\ell_{\text{JJ}}$ is the length of the Josephson junction and $p_{\text{tunnel}}$ is the per-QP tunneling probability at the junction. Combining these definitions, we can write Eq. 63 as
@@
where we have expanded the hyperbolic cotangent term, because $1/\lambda_{d}$ can be made much less than 1 with reasonable design parameters.
For illustrative purposes of our model, we assume AlMn for our trapping material. Since the relative amount of Mn in the AlMn should be very low, the QP scattering diffusion length should be close to that of a typical Al thin film. As such, we use a value of $\alpha_{\text{trap}}\approx 570$ as measured in [73]. From [78] we estimate the QP tunneling probability to be $p_{tunnel}\approx 1-10\,(\times 10^{-6})$ for pure Al. For the AlMn in this model, we adopt the conservative lower bound of $1\times 10^{-6}$.